# From 100 kpc to 10 Gpc: Dark Matter self-interactions before and after DESI

### v1

@@ Abstract
We consider Dark Matter self-interactions mediated by ultralight scalars. We show that effectively massless mediators lead to an enhancement of the matter power spectrum, while heavier mediators lead to a suppression, together with a feature around their Jeans scale. We derive the strongest present constraints by combining Planck and BOSS data. The recent DESI measurements of Baryon Acoustic Oscillations exhibit a mild preference for long-range self-interactions, as strong as 4 per mille of the gravitational coupling. Forthcoming data from DESI itself and Euclid will confirm or disprove such a hint.
@@ I Introduction
Cosmological data on different scales provide evidence for the existence of Dark Matter (DM) and Dark Energy, the first in the form of a collisionless non-relativistic fluid, and the second as a non-zero cosmological constant $\Lambda$ in its simplest incarnation. These two fluids, together with the Standard Model (SM) of particle physics, form the current standard description of the post-inflationary Universe, widely known as $\Lambda$CDM.
The existence of the DM fluid is one of the most definitive indications for the existence of physics beyond the SM, and its characterization is one of the most important challenges of particle physics and cosmology. In perfect analogy with the observed complex structure of the visible sector, it is likely that what we currently view as DM hides a rich Dark Sector (DS), whose dynamics simplifies to the one of a collisionless non-relativistic fluid on cosmological scales. Observing a deviation from $\Lambda$CDM could then provide a hint of the hidden microphysics of the DS.
Thanks to the increasing precision of the available datasets, cosmology is in a unique position to explore the nature of the DS, by testing its properties directly and independently of any portals connecting it to the SM. The exquisite knowledge of the Cosmic Microwave Background (CMB) provided by Planck [1] will soon be accompanied by a comparable precision in measurements of galaxy clustering through the forthcoming results of the DESI survey first [2, 3], the Euclid space telescope shortly thereafter [4], and new observational facilities proposed for the next decade [5, 6, 7, 8].
Due to the need to describe perturbations beyond the linear regime, extracting full information from galaxy clustering data requires a much more advanced modelling compared to the CMB, which is provided by the Effective Field Theory of Large Scale Structure (EFTofLSS) [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Its application to present data has been extremely successful [22, 23, 24, 25, 26, 27, 28] and future data will enable measurements of $\Lambda$CDM parameters to unprecedented precision [29, 30], with the possibilities to test scenarios beyond $\Lambda$CDM largely unexplored.
In this Letter we take a major step forward in understanding the impact of DM self-interactions on cosmological scales. For concreteness, we consider self-interactions mediated by an ultralight scalar field $\varphi$ with trilinear coupling to the DM $\chi$, $\mathcal{L}_{\text{int}}=-\,\kappa\varphi\chi^{2}$ [31, 32, 33, 34, 35]. We expect this simple model to capture most of the phenomenology of attractive DM self-interactions in cosmological observables. Introducing the dimensionless field $s=G_{s}^{1/2}\varphi$, with $G_{s}\equiv\kappa^{2}/m_{\chi}^{4}\equiv 4\pi G_{N}\beta$, we can write at the leading order in a $1/G_{s}$ expansion
@@
which shows that the interaction can be traded for a DM space-time dependent mass, $m_{\chi}(s)=m_{\chi}(1+2s)^{1/2}$. The DM self-interaction is thus characterized by two parameters: its strength normalized to the one of gravity, $\beta$, and its range set by the Compton wavelength of the scalar, $\lambda_{\varphi}=1/m_{\varphi}$. As we show in Sec. II, the latter is responsible for the modifications to the cosmological background. The imprints on the perturbations are instead controlled by the de Broglie wavelength of the scalar moving with the Hubble flow (i.e. the Jeans scale) [36], thus allowing cosmological observables to probe shorter ranges than naively expected.
@@
We focus on Compton wavelengths larger than the size of the horizon at matter-radiation equality, $\lambda_{\varphi}\gtrsim H_{\rm{eq}}^{-1}\sim 30\;\mathrm{kpc}$, for which the EFTofLSS modelling developed in Ref. [37] is expected to hold. We show that only long-range forces with $\lambda_{\varphi}\geq H_{0}^{-1}\sim 4\,\mathrm{Gpc}$ lead to a scale-independent increase in the matter power spectrum compared to $\Lambda$CDM. For $\lambda_{\varphi}<H_{0}^{-1}$ the power spectrum is instead suppressed at any scale, with a step-like feature appearing around the Jeans scale of the massive mediator. An analytical explanation of these results is given in Sec. II.
@@
We obtain the current best limits on the coupling strength of the new force by combining the CMB with BOSS measurements of the full shape of the galaxy power spectrum [38] and previous data on Baryon Acoustic Oscillations (BAO) [39, 40, 41]. [^1] We have checked that replacing BOSS full-shape with BOSS BAO data gives nearly identical results, due to the limited BOSS statistics. This was already observed in Ref. [35] for massless mediators. These limits are shown as a gray solid line in Fig. 1. We then derive the projected sensitivities of the full-shape datasets of DESI, Euclid, and the proposed next-generation galaxy survey MegaMapper (colored lines in Fig. 1). The details are presented in Sec. III.
Finally, we show that the recent DESI BAO measurement [42] seems to mildly favor long-range DM self-interactions over $\Lambda$CDM. The best fit points are shown in red in Fig. 1 and the physics driving this hint is discussed in Sec. IV. As Fig. 1 manifests, this intriguing preference will be conclusively confirmed or disproved by forthcoming full-shape data.
@@
Figure 1: Parameter space of DM self-interactions: $\beta$ is the interaction strength relative to gravity and $(m_{\varphi}/H_{0})^{-1}$ sets the range of the interaction. The dashed gray line is the $95\%$ CL constraint from Planck data [1], the solid gray line is the combination with BOSS full-shape data [38] and previous BAO measurements [39, 40, 41]. The red dots indicate the best fit points of the DESI BAO data [42] with a $68\%$ CL error band. The orange line shows the sensitivity of future DESI full-shape data [2, 3], while the green line corresponds to the forecasted reach of Euclid [43]. The blue line gives the sensitivity of the proposed MegaMapper [5, 6].
@@ II Massless vs massive mediators
For a massive mediator with $H_{0}<m_{\varphi}\lesssim H_{\rm eq}\,$, the mass and DM source terms in Eq. 2 become comparable at some time during matter domination, $a_{m_{\varphi}}\simeq[3H_{\rm eq}/(4\sqrt{2}\,m_{\varphi})]^{2/3}a_{\rm eq}$. From this time on, the quadratic potential quickly dominates and the scalar begins to oscillate around the origin, eventually settling there. Its energy density redshifts like pressureless matter [44] and makes up a fraction of the matter density today that can be estimated as
@@
The first modification to the CDM behavior corresponds to the scale-independent enhancement of growth active when $a<a_{m_{\varphi}}$, while the last two corrections embody the suppressed clustering of $s$ when $a>a_{m_{\varphi}}$. The $-f_{s}$ term simply manifests that a fraction $f_{s}$ of matter does not cluster, whereas the log-enhanced $f_{s}$ term encodes the impact on the $\chi$-baryon perturbations of the suppression of the gravitational potential by a factor $1-f_{s}$. These effects are formally analogous to those of free-streaming massive neutrinos [46], although the underlying physical phenomena are distinct. For a massive mediator the parametrically large $f_{s}$ (see Eq. 3) gives the dominant effect in Eq. 4, causing an overall suppression of the linear power spectrum compared to $\Lambda$CDM and a step-like feature at $k\sim k_{\rm J}(a)$. Conversely, for a massless mediator one finds a nearly scale-independent increase of power, as previously shown in Refs. [35, 37].
In this Letter we focus on masses $m_{\varphi}\lesssim H_{\rm eq}$, which satisfy $k_{\rm J}(a)\lesssim k_{\rm eq}\equiv a_{\rm eq}H_{\rm eq}$. This ensures that $k\gtrsim k_{\rm J}$ for the modes of relevance to LSS, namely $k_{\rm eq}\lesssim k\lesssim k_{\rm NL}$ (where $k_{\rm NL}$ is the nonlinear scale). Therefore the EFTofLSS as derived for a massless mediator [37] still applies. Nonetheless, the theoretical uncertainties in our non-linear modelling increase toward the upper end of the considered mass window, $m_{\varphi}/H_{0}\gtrsim 10^{4}$, because (i) eventually $k_{\rm J}$ becomes larger than $k_{\rm eq}$, hence a more careful treatment of the behavior of a relativistic fluid in the EFTofLSS is required; and (ii) the time logarithm $\log a_{m_{\varphi}}/a_{\rm eq}$ appearing in Eq. 4 is reduced, hence non-log-enhanced $\mathcal{O}(\beta)$ corrections become more important (see Appendix A for further details).
@@ III Results
Current and projected bounds in the 2-dimensional parameter space of $(H_{0},\beta)$ are presented in Fig. 2. The left panel shows the case $m_{\varphi}/H_{0}=0.1$, while the right one corresponds to $m_{\varphi}/H_{0}=100$. Different colors indicate the combinations of present and future LSS datasets with Planck CMB measurements. [^2] The forecasts have been obtained using the FishLSS code [29, 47]. We use the default samples with the exception of DESI, where we add the Bright Galaxy Survey to the default Emission Line Galaxies (ELG). The galaxy bias expansion is written in terms of the $\chi$-baryon field, as the scalar field does not cluster on the relevant scales [48, 49]. To obtain the forecasts we first invert the covariance matrix of the Markov Chain Monte Carlo (MCMC) chains from Planck+BOSS data to estimate the Fisher matrix. Next, we add it to the Fisher matrix of the LSS forecasts obtained with FishLSS. We then generate a Gaussian sample centered at the MCMC best fit value (setting $\beta$ to 0) with covariance given by the inverse of the combined Fisher matrix. Finally we drop all points with $\beta<0$ and obtain the 68% and 95% CL intervals.
@@ V Outlook
In this Letter we showed how galaxy clustering data provides valuable information about DM self-interactions independently of any couplings between DM and the visible sector. We derived world-leading cosmological constraints on attractive self-interactions with range $\lambda_{\varphi}\gtrsim H_{\text{eq}}^{-1}\sim 30$ kpc, based on a combination of the full shape of the BOSS galaxy power spectrum with Planck CMB data and BAO measurements. Forthcoming galaxy clustering data will improve the sensitivity by around one order of magnitude, exploring a vast and uncharted parameter space of DM self-interactions.
The EFTofLSS allows us to consistently account for the new physics parameters together with the parameters encoding our partial knowledge about the clustering (bias coefficients and ultraviolet counterterms, see Appendix A). Fitting all these parameters simultaneously to data makes the derived constraints robust against theoretical systematics. For the ranges considered here, precision cosmology outperforms complementary astrophysical probes [57, 58, 59].
Remarkably, the distinctive features of massive mediators make the power spectrum able to distinguish self-interactions of different ranges. As an example of this discriminating potential, we show that the recent Year-1 DESI BAO measurements signal a mild preference for a long-range self-interaction, which would accommodate the higher value of $H_{0}$ observed by DESI. Future DESI full-shape data, as well as Euclid data, will be able to fully confirm or disprove such a hint.
@@ Appendix A Massive vs massless forces in detail
where we defined the fraction of the total energy density in the species $i$ as $\Omega_{i}\equiv\bar{\rho}_{i}/\bar{\rho}_{\rm tot}$.
@@
where $f_{s}^{\rm massless}=\beta\widetilde{m}_{s}^{2}f_{\chi}^{2}/3$ is the result one finds [35] for $m_{\varphi}\leq H_{0}\,$, by employing the solution in Eq. 6. When providing analytical results in the main text, such as Eq. 3, we have taken $\mathcal{G}=1$ as representative value.

### v2

@@ Abstract
We consider Dark Matter self-interactions mediated by ultralight scalars. We show that effectively massless mediators lead to an enhancement of the matter power spectrum, while heavier mediators lead to a suppression, together with a feature around their Jeans scale. We derive the strongest present constraints by combining Planck and BOSS data. The recent DESI measurements of Baryon Acoustic Oscillations exhibit a mild $2\hskip 0.56905pt\sigma$ preference for long-range self-interactions, as strong as 4 per mille of the gravitational coupling. Full-shape analyses of forthcoming DESI and Euclid data will confirm or disprove such a hint.
@@ I Introduction
Cosmological observations on different scales have provided compelling evidence for the existence of Dark Matter (DM) and Dark Energy (DE), the first in the form of a collisionless non-relativistic fluid, and the second as a non-zero cosmological constant $\Lambda$ in its simplest incarnation. Together with the Standard Model (SM) of particle physics, these two fluids form the current standard description of the post-inflationary Universe, widely known as $\Lambda$CDM.
The existence of the DM fluid is one of the most definitive indications of phenomena beyond the SM, making its characterization a central challenge in both particle physics and cosmology. In perfect analogy with the observed complex structure of the visible sector, it is likely that what we currently view as DM hides a rich Dark Sector (DS), whose dynamics simplifies to the one of a collisionless non-relativistic fluid on cosmological scales. Observing a deviation from $\Lambda$CDM could then provide crucial insights on the hidden microphysics of the DS.
As observational precision improves, cosmology is uniquely positioned to test the properties of the DS directly, independent of any interactions connecting it to the SM. The exquisite knowledge of the Cosmic Microwave Background (CMB) provided by Planck [1] will soon be complemented by a comparable precision in measurements of galaxy clustering by the DESI survey first [2, 3], the Euclid space telescope shortly thereafter [4], and new observational facilities proposed for the next decade [5, 6, 7, 8].
Due to the need to describe perturbations beyond the linear regime, extracting full information from galaxy clustering data requires an advanced modelling provided by the Effective Field Theory of Large Scale Structure (EFTofLSS) [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. The EFTofLSS has been successfully applied to present data [22, 23, 24, 25, 26, 27, 28] and is expected to provide unprecedented accuracy in the measurement of $\Lambda$CDM parameters when applied to future datasets [29, 30]. In this article, we extend the EFTofLSS framework to the exploration of DM dynamics beyond the minimal $\Lambda$CDM model, focusing on the implications of DM self-interactions on cosmological scales.
For concreteness, we consider self-interactions mediated by an ultralight scalar field $\varphi$ with trilinear coupling to the DM $\chi$, $\mathcal{L}_{\text{int}}=-\,\kappa\varphi\chi^{2}$ [31, 32, 33, 34, 35]. We expect this simple model to capture most of the phenomenology of attractive DM self-interactions in cosmological observables. Introducing the dimensionless field $s=G_{s}^{1/2}\varphi$, with $G_{s}\equiv\kappa^{2}/m_{\chi}^{4}\equiv 4\pi G_{N}\beta$, we can write at the leading order in a $1/G_{s}$ expansion
@@
which shows that the interaction can be traded for a DM space-time dependent mass, $m_{\chi}(s)=m_{\chi}(1+2s)^{1/2}$. The DM self-interaction is characterized by two parameters: its strength normalized to the one of gravity, $\beta$, and its range set by the Compton wavelength of the scalar, $\lambda_{\varphi}=1/m_{\varphi}$. As we show in Sec. II, the latter is responsible for the modifications to the cosmological background. The imprints on the perturbations are instead controlled by the de Broglie wavelength of the scalar moving with the Hubble flow (i.e. the Jeans scale) [36], thus allowing cosmological observables to probe shorter ranges than naively expected.
@@
We focus on mediators with de Broglie wavelengths larger than the comoving horizon at matter-radiation equality, or equivalently Jeans wavenumbers $k_{\rm J}\lesssim 0.01\,h\,\mathrm{Mpc}^{-1}$, for which the EFTofLSS modelling developed in Ref. [37] is expected to hold. This implies that we can probe mediator masses as large as $m_{\varphi}\sim 10^{5}\,H_{0}\approx 10^{-28}\;\mathrm{eV}$. We show that only long-range forces with $\lambda_{\varphi}\geq H_{0}^{-1}\sim 4\,\mathrm{Gpc}$ lead to a scale-independent increase in the matter power spectrum compared to $\Lambda$CDM. For $\lambda_{\varphi}<H_{0}^{-1}$ the power spectrum is instead suppressed at any scale, with a step-like feature appearing around the Jeans scale of the massive mediator. An analytical explanation of these results is given in Sec. II.
@@
We obtain the current best limits on the coupling strength of the new force by combining the CMB with BOSS measurements [38] of the full shape of the galaxy power spectrum, which we calculate using the EFTofLSS, and previous data on Baryon Acoustic Oscillations (BAO) [39, 40, 41]. [^1] Replacing BOSS full-shape with BOSS BAO data gives nearly identical results, due to the limited BOSS statistics. This was already observed in Ref. [35] for massless mediators. These limits are shown as a gray solid line in Fig. 1. For each value of $m_{\varphi}/H_{0}$, the interaction strength $\beta$ is scanned in a Markov Chain Monte Carlo (MCMC) analysis, simultaneously with the standard cosmological parameters and the additional parameters of the EFTofLSS. We then derive the projected sensitivities of the full-shape datasets of DESI, Euclid, and the proposed next-generation galaxy survey MegaMapper (colored lines in Fig. 1). The details are presented in Sec. III. The constraints and forecasts we report in Fig. 1 are the strongest to date in the whole window of DM self-interaction ranges considered.
@@
Figure 1: Parameter space of DM self-interactions: $\beta$ is the interaction strength relative to gravity and $(m_{\varphi}/H_{0})^{-1}$ sets the range of the interaction. The dashed gray line is the $95\%$ CL constraint from Planck data [1], the solid gray line is the combination with BOSS full-shape data [38] and previous BAO measurements [39, 40, 41]. The red dots indicate the best fit points of the DESI BAO data [42] with a $95\%$ CL error band (the $68\%$ CL error band is shown in darker red). The orange line shows the sensitivity of future DESI full-shape data [2, 3], while the green line corresponds to the forecasted reach of Euclid [43]. The blue line gives the sensitivity of the proposed MegaMapper [5, 6].
Finally, we show that the recent DESI BAO measurement [42] seems to mildly favor long-range DM self-interactions over $\Lambda$CDM. The best fit points are shown in red in Fig. 1 and the physics driving this hint is discussed in Sec. IV. This intriguing preference will be conclusively confirmed or disproved by forthcoming full-shape data.
@@ II Massless vs massive mediators
For a massive mediator with $H_{0}<m_{\varphi}\lesssim H_{\rm eq}\,$, the mass and DM source terms in Eq. 2 become comparable at some time during matter domination, $a_{m_{\varphi}}\simeq[3H_{\rm eq}/(4\sqrt{2}\,m_{\varphi})]^{2/3}a_{\rm eq}$. From this time on, the quadratic potential quickly dominates and the scalar begins to oscillate around the origin, eventually settling there. Its energy density redshifts like pressureless matter [44] and makes up a fraction of the matter density today that can be estimated as ($\log^{2}x\equiv(\log x)^{2}$)
@@
The first modification to the CDM behavior corresponds to the scale-independent enhancement of growth active when $a<a_{m_{\varphi}}$, while the last two corrections embody the suppressed clustering of $s$ when $a>a_{m_{\varphi}}$. The $-f_{s}$ term simply manifests that a fraction $f_{s}$ of matter does not cluster, whereas the log-enhanced $f_{s}$ term encodes the impact on the $\chi$ plus baryon perturbations of the suppression of the gravitational potential by a factor $1-f_{s}$. These effects are formally analogous to those of free-streaming massive neutrinos [46], although the underlying physical phenomena are distinct. For a massive mediator, the parametrically large $f_{s}$ (see Eq. 3) gives the dominant effect in Eq. 4, causing an overall suppression of the linear power spectrum compared to $\Lambda$CDM and a step-like feature at $k\sim k_{\rm J}(a)$. Conversely, for a massless mediator one finds a nearly scale-independent increase of power, as previously shown in Refs. [35, 37].
In this article we focus on masses $m_{\varphi}\lesssim H_{\rm eq}$, corresponding to $\lambda_{\varphi}\gtrsim 30\;\mathrm{kpc}$, which satisfy $k_{\rm J}(a)\lesssim k_{\rm eq}\equiv a_{\rm eq}H_{\rm eq}$. This ensures that $k\gtrsim k_{\rm J}$ for the modes of relevance to LSS, namely $k_{\rm eq}\lesssim k<k_{\rm NL}$ (where $k_{\rm NL}$ is the nonlinear scale). Therefore the EFTofLSS as derived for a massless mediator [37] still applies. Nonetheless, the theoretical uncertainties in our non-linear modelling increase toward the upper end of the considered mass window, $m_{\varphi}/H_{0}\gtrsim 10^{4}$, because (i) eventually $k_{\rm J}$ becomes larger than $k_{\rm eq}$, hence a more careful treatment of the behavior of a relativistic fluid in the EFTofLSS is required; and (ii) the time logarithm $\log a_{m_{\varphi}}/a_{\rm eq}$ appearing in Eq. 4 is reduced, hence non-log-enhanced $\mathcal{O}(\beta)$ corrections become more important (see Appendix A for further details).
@@ III Results
Current and projected bounds in the 2-dimensional parameter space of $(H_{0},\beta)$ are presented in Fig. 2. The left panel shows the case $m_{\varphi}/H_{0}=0.1$, while the right one corresponds to $m_{\varphi}/H_{0}=100$. Different colors indicate the combinations of present and future LSS datasets with Planck CMB measurements. The forecasts have been obtained using the FishLSS code [29, 47], see Appendix A for additional details.
@@ V Outlook
In this article, we showed how galaxy clustering data reveals powerful insights on DM self-interactions, independent of any couplings with the visible sector. We presented world-leading cosmological constraints on attractive self-interactions with range $\lambda_{\varphi}\gtrsim H_{\text{eq}}^{-1}\sim 30$ kpc, derived from our analysis of the BOSS galaxy power spectrum within the EFTofLSS, combined with Planck CMB data and BAO measurements. Forthcoming galaxy clustering data will improve the sensitivity by around one order of magnitude, exploring a vast and largely uncharted parameter space of DM self-interactions.
The EFTofLSS allows us to consistently incorporate new physics parameters alongside those that encode our partial understanding of the clustering (bias coefficients and ultraviolet counterterms, see Appendix B). By fitting all these parameters simultaneously to data, we enhance the robustness of the derived constraints against theoretical systematics. For the ranges considered here, precision cosmology outperforms complementary astrophysical probes, establishing itself as a powerful tool for understanding DM self-interactions [55, 56, 57].
Remarkably, the distinctive features of massive mediators enable the power spectrum to differentiate between self-interactions of different ranges. As an example of this discriminating potential, we showed that the recent Year-1 DESI BAO measurements signal a mild preference for long-range self-interactions, which would accommodate the higher value of $H_{0}$ observed by DESI. Future full-shape data from DESI, along with data from Euclid, will be able to fully confirm or disprove such a hint.
@@ Appendix A Massive vs massless forces in detail
where we defined $\Omega_{i}\equiv\bar{\rho}_{i}/\bar{\rho}_{\rm cr}$ for the $i$ species, with $\bar{\rho}_{\rm cr}=3H_{0}^{2}/(8\pi G_{N})$.
@@
where $f_{s}^{\rm massless}=\beta\widetilde{m}_{s}^{2}f_{\chi}^{2}/3$ is the result one finds [35] for $m_{\varphi}\leq H_{0}\,$, by employing the solution in Eq. 6. When providing analytical results in the main text, such as Eq. 3, we have taken $\mathcal{G}=1$ as representative value. Notice that in this work and in Refs. [35, 37] we defined $\Omega_{i}=\bar{\rho}_{i}/\bar{\rho}_{\rm cr}$ for the individual species but $\Omega_{m}=\bar{\rho}_{m}/\bar{\rho}_{\rm tot}(a)$, hence $\Omega_{\chi}+\Omega_{b}+\Omega_{s}\neq\Omega_{m}$ in Eq. 10.
@@
Lastly, we provide some technical details about our forecasts, which have been derived using FishLSS [29, 47]. We take the default samples with the exception of DESI, where we add the Bright Galaxy Survey to the default Emission Line Galaxies (ELG). As already explained, the galaxy bias expansion is written in terms of the $\chi$-baryon field, as the scalar field does not cluster on the relevant scales [69, 68]. To obtain the forecasts we first invert the covariance matrix of the MCMC chains from Planck+BOSS data to estimate the Fisher matrix. Next, we add it to the Fisher matrix of the forecasts obtained with FishLSS. We then generate a Gaussian sample centered at the MCMC best fit value (setting $\beta$ to 0) with covariance given by the inverse of the combined Fisher matrix. Finally, we drop all points with $\beta<0$ and obtain the 68% and 95% CL intervals.